# Collapse of Irrelevant Representations (CIR) Ensures Robust and Non-Disruptive LLM Unlearning

### v1

@@ Abstract
Current unlearning techniques and safety training consistently fail to remove dangerous knowledge from language models. We analyze the root causes and propose a highly selective technique which unlearns robustly and without disrupting general performance.
We perform PCA on activations and module output gradients to identify subspaces containing common representations, and collapse them before calculating unlearning updates. This way we avoid unlearning general representations, and only target those specific to the unlearned facts.
When unlearning WMDP dataset facts from Llama-3.1-8B, we drop post-attack accuracy 80× more than our best baseline (Circuit Breakers) on biohazardous facts and 30× more on cyberhazardous facts. Despite this, we disrupt general performance 30× less (only 0.1% WikiText loss increase), while requiring less than 3 GPU-seconds per fact. Code: github.com/filyp/unlearning
@@ 1 Introduction
During pre-training, language models learn hazardous capabilities useful e.g. for bioterrorism and cybercrime (Li et al., 2024). They even acquire information about their own safety controls, which in the future could let models circumvent them (Roger, 2024; Greenblatt et al., 2024).
Popular safety training approaches like DPO and RLHF do not eliminate unwanted capabilities, but rather teach the model to stop using them Lee et al. (2024). These concealed capabilities can be resurfaced by jailbreak attacks (Zou et al., 2023) or even completely accidentally (Qi et al., 2023). Even methods designed specifically for unlearning can be easily reversed (Łucki et al., 2025; Lynch et al., 2024; Deeb & Roger, 2024).
In this work, we identify the fundamental cause of unlearning failure: naive unlearning disrupts general representations shared between harmful and benign capabilities (see Section 3.3). Then, during fine-tuning attacks, these broken representations can be identified and fixed because they are also present in the attacker’s training data. We saw that unlearning becomes vulnerable to attacks as soon as it causes even 0.1% general performance degradation.
Figure 1 presents the CIR technique, which before calculating unlearning updates first removes the general representations from activations and module output gradients. [^1] Disambiguation: By ”gradients” we always mean the module output gradients, ones that flow into modules during backpropagation before weight updates are computed. For the final per-weight gradients, we always use the term ”update”. We pair it with a representation engineering loss. But since knowledge is mainly stored in MLPs (Nanda et al., 2023), instead of breaking residual stream activations as in Zou et al. (2024), we propose the MLP breaking loss, which breaks MLP outputs before they are added to the residual stream. This works 40% better.
(a) Collapse of Irrelevant Representations (CIR) diagram.
In orange we show ”dirty” vectors, which contain representations irrelevant to the unlearning task. Unlearning on them would cause disruption and unrobustness. In green we show the purified vectors, which target only the unwanted representations.
(b) Comparison of unlearning methods on WMDP-Bio.
Methods are terminated once they hit a disruption threshold and then tested under a fine-tuning attack. Like Deeb & Roger (2024), during the attack we retrain on facts different than evaluated facts, but from the same category.
@@ 2 Related work
Methods relying solely on backpropagation, like DPO (Rafailov et al., 2024), only deactivate unwanted capabilities, not remove them (Lee et al., 2024). For this reason, alternative unlearning approaches have been proposed. Several recent methods aim to disrupt the intermediate activations of the model (Zou et al., 2024; Rosati et al., 2024; Li et al., 2024). Others incorporate meta-learning (Tamirisa et al., 2024; Sondej et al., 2025; Henderson et al., 2023) which simulates how an attacker could relearn the unwanted knowledge, to prepare against it. Some try to locate the harmful neurons or activation directions and ablate them Wang et al. (2024); Wu et al. (2023); Uppaal et al. (2024); Suau et al. (2024).
@@
However, currently all existing unlearning techniques are easily reversed by fine-tuning, jailbreaks, few-shot prompting, disabling refusal mechanisms, or out-of-distribution inputs (Łucki et al., 2025; Lynch et al., 2024). Even for methods which ablate harmful concepts, Lo et al. (2024) found that the model can repurpose neurons with similar meaning to quickly relearn them.
@@
Failure of current unlearning methods has been shown most explicitly by Deeb & Roger (2024), where attackers could recover supposedly unlearned facts by training on a completely independent set of facts, which definitively proves that they were not removed. In our experiments, for the fine-tuning attacks we use the same approach – trying to recover the target facts by training on different facts from the same category.
3 Motivation for unlearning selectivity
In this section, we will share our insights as to why unlearning has been so challenging. We hope to show how our technique emerges naturally as a response to these issues. (To go straight to our method, you can skip to Section 4.)
@@ 3.1 Disruption leads to unrobustness
Existing unlearning methods are consistently easy to undo. We have noticed that we can predict how successful a fine-tuning attack will be by looking at the disruption during unlearning. Let us divide unlearning runs into two phases: ”non-disruptive”, which lasts as long as retain loss stays below 100.1% of its initial value, and ”disruptive”, which starts after that.
On Figure 5 (in the appendix) we see that unlearning achieved during the disruptive phase is usually reversible with a fine-tuning attack. But surprisingly, unlearning that happened without any disruption remains robust. Sometimes disruptive unlearning is partially robust too, but it is not guaranteed.
This means that letting unlearning disrupt general performance is unacceptable. In our experiments, unlearning can become unrobust after as little as 0.1% retain set disruption. This finding explains the results from Deeb & Roger (2024), who allowed unlearning to disrupt retain loss by 5%, and then showed near-zero robustness.
@@ 3.2 Disruption is costly
Existing unlearning techniques also try to minimize disruption, but they typically do it by training on the retain set, hoping to undo the damage that the unlearning has caused. But while breaking the model is easy, in our experience fixing it takes a prohibitively long time. This makes sense – after all the model’s weights have already been extremely optimized through multi-million dollar training runs. If we aimlessly break them, we should not expect that going back to the optimal values will be easy. So instead of fixing the damage from unlearning, ideally we should not cause the damage in the first place.
@@ 3.3 Disruption of superficially similar facts
Unlearning modifies the model to make unwanted answers less likely. For example when unlearning ”The capital of France is Paris”, there are many ways to make ”Paris” less likely: actually forgetting it is France’s capital, forgetting what ”capital” means, forgetting that ”is” requires the answer to follow, etc. In fact, as Figure 2 shows, unlearning ”The capital of France is Paris”, accidentally unlearns ”The capital of Spain is Madrid84% as strongly. (We unlearn only the tokens shown in purple.) It can even affect completely unrelated facts. Interestingly, wrong facts are not disrupted. See Appendix A for more examples.
Similarly, unlearning biohazardous facts likely disrupts many benign biological concepts. This could explain why we can recover ”unlearned” facts by retraining on unrelated biological text (Deeb & Roger, 2024) – retraining fixes these benign concepts.
Figure 2: Disruption caused by unlearning a simple fact. We show how unlearning ”The capital of France is Paris” disrupts the recall of other facts. We measure disruption using cosine similarity between the model’s update on the ”Paris” fact and the evaluated fact. Activations column shows a slice of activations incoming into a middle layer MLP module on the token position right before the answer. Gradients column shows a slice of the gradients incoming into the same module during backpropagation when aiming to unlearn the answer (in purple).
In Figure 2 we see that the activations (and to a lesser extent, gradients) are quite similar across different facts. This sheds light on why superficially similar facts are disrupted – most representations are not specific to the fact we are trying to unlearn, but more general. And since the model updates are computed using these ”dirty” activations and gradients, other facts which also contain these general representations will be disrupted. To prevent this, we need to find a way to filter out these general representations.
@@ 3.4 Filtering out disruption is easier in activation and gradient space
A natural thing to try if we want to be selective, is to limit which weights to update. For example Sondej et al. (2025) have shown unlearning improvements when allowing to modify only the weights where the signs of the unlearning and the retaining update are the same. Similarly, the A-GEM technique (Chaudhry et al., 2019) from the field of continual learning aims to avoid performance disruption by projecting the weight updates to make them orthogonal to the retaining updates. Such projections have also been successfully used for unlearning (Wu et al., 2025).
On Figure 3, under ”masked per weight” you can see the effects of such filtering techniques. They significantly reduce the disruption (shown in red), but some of it still escapes the filtering. That is because the ”control/retaining updates” that we use to decide which weights to filter out never match the actual disruption perfectly. (Compare the blue control pattern and the red disruption pattern.)
Figure 3: Comparison of two masking strategies.
We show a slice of the updates of one weight matrix when unlearning ”The capital of France is Paris”. We color a weight green if its update successfully transfers to unlearning ”France’s capital is Paris”, and red if it disrupts the recall of ”The capital of Spain is Madrid”.
We also record disruption of a control fact: ”The capital of Italy is Rome” (shown in blue). Then we use this control disruption as a guess to which weights (or columns and rows) are most disruptive, and filter the unlearning update accordingly.
Can we improve this filtering? When we look at update patterns, we see that both disruption and transfer appear as column- and row-wise stripes. (This happens because updates are computed by multiplying the activations with the gradients, which makes the update low-rank.) It looks like it is certain rows and columns that are disruptive, rather than individual weights.
Since the disruption patterns shift within these columns and rows, it means that granular, per-weight filtering will miss some weights. It makes more sense to identify and remove whole faulty columns and rows (which would correspond to ablating values in the activations and module output gradients). Indeed, we see that it improves the disruption/transfer ratio from 33% to 5%.  [^2] Another advantage of intervening on whole columns and rows, is that we can save memory by operating on the activations and module output gradients rather than on the final update.
@@ 4 Collapse of Irrelevant Representations
Following the findings from the previous section, we have tried many ways to remove representations which cause disruption. Simply ablating elements of the activations and gradients (like on Figure 3) often struggles to only get rid of disruption. That is because representations exist in superposition (Elhage et al., 2022), so one element takes part in encoding multiple representations, some relevant to the unlearning task, some not.
@@
We found that rather than ablating, it is much better to project out irrelevant representations. Trying to define irrelevant representations manually would be prohibitively tedious, so we assume that if a representation is commonly present in many training texts, it is probably irrelevant. Removing them will leave us only with representations which are specific to the given training text. The most natural way to locate the subspace with most common representations, is to first take the mean of the values and then also their principal components (PCs). (To simplify, we treat the mean as the ”0th principal component”, and whenever we talk about collapsing components, we first collapse the mean.) Equation 1 shows how to collapse activation PCs, and the same is done for gradients.
@@
Based on grid searches shown in Figure 7(a) & 7(b) we chose to project 24 activation PCs and 36 gradient PCs. Performance plateaus for values between 12 and 48 (for both), making precise tuning unnecessary in this range. Removing the first few activation PCs is crucial.
@@
We calculate PCs for each trained MLP module, both for their incoming activations and for their module output gradients incoming during backpropagation. Normally these activations and module output gradients would be Einstein summed to calculate the weight updates. But we discard this default weight update and instead first collapse PCs and only then calculate the update. PCs drift over time, so it helps to recompute them during unlearning – we do it after each epoch, but it can be rarer. PCs can be computed over any dataset, but we have found it works best to simply compute them on the unlearning corpus itself. This luckily makes the algorithm much more efficient, because we can reuse forward and backward passes for unlearning and for fetching activations and gradients. See Algorithm 1 for the pseudocode.
We only intervene on MLPs, since this is where the model’s knowledge is stored (Nanda et al., 2023). Also, collapsing representations on attention modules would be complex and specific to the model implementation.
@@
CIR is compatible with any unlearning loss function and (optionally) with any retain loss function. First we tried loss functions which operate on the final logits, like negative cross entropy, negative entropy (Tamirisa et al., 2024), or (best in this category) simply minimizing the logit for the target token (but not below 0). The last one is extremely good at preventing the model from recalling the harmful answer, but does not generalize to preventing recognizing the harmful answer during multiple-choice questions. So it may be the optimal choice if we only care for recall, not recognition, but it must be checked whether it fails to generalize in some other ways.
@@
In contrast, losses which aim to break intermediate representations prevent both recall and recognition of the harmful answers. The SOTA representation breaking method is Circuit Breakers (Zou et al., 2024), which minimizes (but only down to 0) the cosine similarity between current and initial activations of the residual stream.
We improve on this SOTA in two ways. First, we notice a problem with cosine similarity: it can be minimized not only by removing the original representation, but also by adding some big random direction. We expected this to be disruptive, so we replaced cosine similarity with the dot product. Indeed, on Figure 6 the dot product disrupts the model much less for the same amount of unlearning, and we see that it is connected to cosine similarity growing the activation norm.
Secondly, rather than breaking activations on the residual stream (which contains representations added there by both MLPs and attention layers), we decided to work at the source, and directly break the MLP outputs before they are added to the stream. Figure 7(c) shows that this improves unlearning vs disruption by an additional 40%, and that it is best to target MLPs on layers 6-12 (for a 32-layer Llama 8B). [^3] It also means that it is enough to do forward and backward passes on just the first 12 layers, which is a major speedup. So our final unlearning loss is:
@@
We normalize with the average norm of the original MLP outputs, because later layers have bigger norms and could dominate the loss too much. Lastly, we also decided not to break representations following the <BOS> token, as this disrupts all texts, including benign ones. We also train on the retain set, using the representation engineering loss: $\text{resid\_stream}_{act}-\text{resid\_stream}_{orig\_act}||$ from the original circuit breakers paper (Zou et al., 2024), which penalizes changing residual stream activations on the retain set.
Figure 4: WMDP-Cyber unlearning results.
Circuit Breakers had an abrupt unlearning reversal where apparently its retain loss component started undoing the gains, so we have run a second relearning run from the point of minimum accuracy, but it turned out even less robust. We also run another CIR run with a higher allowed disruption (1%, because 3% can be hard to reach for CIR).
5 Method comparisons

### v2

@@ Abstract
Current unlearning and safety training methods consistently fail to remove dangerous knowledge from language models. We identify the root cause – unlearning targets representations which are too general – and develop a highly selective technique that unlearns robustly while preserving general performance.
Our method performs PCA on activations and module-output gradients to identify subspaces containing common representations, then collapses these subspaces before computing unlearning updates, a technique we term Collapse of Irrelevant Representations (CIR). This avoids unlearning general knowledge and targets only representations specific to the facts being unlearned.
When unlearning bio- and cyber-hazardous facts from Llama-3.1-8B, we achieve over 30× greater reduction in post-attack accuracy than the best baseline (Circuit Breakers), while disrupting general performance 30× less, and using less than 3 GPU-seconds per fact.
Thus, by disentangling harmful and benign capabilities at the level of representations, CIR enables robust and non-disruptive unlearning. Our code is available at: github.com/filyp/unlearning
@@ 1 Introduction
During pre-training, large language models (LLM) learn hazardous capabilities useful for bioterrorism and cybercrime (li_wmdp_2024). They even acquire information about their own safety controls, which could enable future models to circumvent them (roger_case_2024; greenblatt_alignment_2024).
Popular safety training approaches (RLHF, DPO) do not eliminate unwanted capabilities, but rather teach the models to stop using them (lee_mechanistic_2024). These concealed capabilities can be resurfaced via jailbreak attacks (zou_universal_2023) or even accidentally through benign fine-tuning (qi_fine-tuning_2023). Moreover, even methods designed specifically for unlearning have been found to be easily reversible through fine-tuning attacks and other adversarial methods (lucki_adversarial_2025; lynch_eight_2024; deeb_unlearning_2024).
In this work, we identify the fundamental cause of unlearning failure: naive unlearning disrupts general representations shared between harmful and benign capabilities (see Section 3.3). Therefore, during fine-tuning attacks, these broken representations can be identified and repaired because they are also present in the attacker’s training data. As evidence, we observe that unlearning becomes vulnerable to attacks as soon as it induces even 0.1% general performance degradation (Section 3.1). This explains the near-zero robustness observed e.g. by deeb_unlearning_2024 when allowing for much higher performance degradation.
To address this issue, we propose a novel technique called Collapse of Irrelevant Representations (CIR). Figure 1 presents the CIR technique, which removes the general representations from activations and module-output gradients before calculating unlearning updates. [^1] Note: By “gradients” we always mean the module-output gradients that flow into modules during backpropagation before weight updates are computed. For the final per-weight gradients, we always use the term “update”. We pair it with a representation engineering loss (zou_improving_2024), which aims to make internal representations orthogonal to the original representations. Prior work targeted representations in the residual stream, but since LLM knowledge is mainly stored in MLP modules (nanda2023factfinding), we propose the MLP breaking loss that instead directly targets MLP outputs before they are added to the residual stream, which improves unlearning selectivity by 40%.
(a) Collapse of Irrelevant Representations (CIR) pipeline. The orange boxes show the “dirty” vectors, which contain representations irrelevant to the unlearning target (see Section 3.3). Unlearning on them would cause disruption and poor robustness. The green boxes show the collapsed vectors (“purified”), which target only the unwanted representations.
(b) Comparison of unlearning methods on WMDP-Bio (li_wmdp_2024) Methods are terminated once they hit a disruption threshold and then tested under a fine-tuning attack. Following deeb_unlearning_2024, during the attack we retrain on facts different than evaluated facts, but from the same category. CIR reaches near ideal robustness to relearning, despite disrupting the WikiText loss 30× less that the baselines.
@@ 2 Related work
Unlearning aims to remove dangerous knowledge and capabilities from LLMs. Methods relying solely on backpropagation, such as DPO (rafailov_direct_2024), only deactivate unwanted capabilities, not remove them (lee_mechanistic_2024). For this reason, alternative unlearning approaches have been proposed. Several recent methods aim to disrupt the intermediate activations of models (zou_improving_2024; rosati_representation_2024; li_wmdp_2024). Others incorporate meta-learning (tamirisa_tamper-resistant_2024; sondej_robust_2025; henderson_self-destructing_2023) which simulates how an attacker could relearn the unwanted knowledge to prepare against it. Some try to locate the harmful neurons or activation directions and then ablate them wang_large_2024; wu_depn_2023; uppaal_detox_2024; suau_whispering_2024.
@@
However, currently all existing unlearning techniques are easily reversed by fine-tuning, jailbreaks, few-shot prompting, disabling refusal mechanisms, or out-of-distribution inputs (lucki_adversarial_2025; lynch_eight_2024). Even for methods which ablate harmful neurons, lo_large_2024 found that the model can repurpose neurons with similar meaning to quickly relearn them.
@@
Failure of current unlearning methods has been shown most explicitly by deeb_unlearning_2024, where attackers could recover supposedly unlearned facts by training on a completely independent set of facts, which clearly shows that they were not removed. Our fine-tuning attacks use the same approach: we try to recover the target facts by training on different facts from the same category. Such attacks do not assume that the attacker has full knowledge of the unlearning dataset, which would be unrealistic.
3 Identifying problems with unlearning
In this section, we share our insights on why unlearning has been so challenging. We hope to show how our technique emerges naturally as a response to these issues. To go straight to our method, skip to Section 4.
@@ 3.1 Disruption leads to unrobustness
Figure 2: Success of fine-tuning attacks is determined by disruption during unlearning. We show 50 unlearning runs, each followed by the same fine-tuning attack. (Details in Appendix E.) For each run, we mark on the y axis the WMDP accuracy that was reached with minimal disruption (less than 0.1%), and we continue unlearning after this 0.1% threshold. During the attack, WMDP accuracy is partially restored (see the arrows), but at most to its level at the disruption threshold (shown in red). It means that only unlearning that happened after the disruption threshold can be reverted, and unlearning that happened without disruption remains robust.
Existing unlearning methods are consistently easy to undo. We observe that the success of a fine-tuning attack can be predicted from the disruption caused during unlearning. To formalize this, we divide unlearning runs into two phases: non-disruptive, which lasts as long as retain loss stays below 100.1% of its initial value, [^2] We found by trial and error that this is the highest disruption threshold for which the robustness guarantee shown on Figure 2 holds. and disruptive, which begins once this threshold is exceeded. (Retain loss is the model’s loss computed over the retain datasets defined in Section 5.)
On Figure 2, we see that unlearning achieved in the disruptive phase is usually reversible by a fine-tuning attack. In contrast, unlearning that occurs without any disruption remains robust.
This shows that allowing unlearning to disrupt general performance is unacceptable. In our experiments, unlearning robustness can collapse after as little as 0.1% retain set disruption. This finding explains the results of deeb_unlearning_2024, who permitted a 5% disruption of the retain loss and observed near-zero robustness.
@@ 3.2 Disruption is costly
Existing unlearning methods also aim to minimize disruption, typically by retraining on a retain set to undo the damage (zou_improving_2024; rosati_representation_2024). While breaking a model is easy, in our experience, repairing it is prohibitively time-consuming and costly because the weights are already finely tuned by large-scale pre-training. Therefore, rather than relying on expensive post-hoc fixes, we should design unlearning methods that avoid causing damage in the first place.
@@ 3.3 Disruption of superficially similar facts
Unlearning modifies the model to make unwanted answers less likely. For example when unlearning “The capital of France is Paris”, there are many ways to make ”Paris” less likely: actually forgetting that it is France’s capital, forgetting what “capital” means, or forgetting that the word “is” requires the answer to follow, etc. In fact, as Figure 3 shows, unlearning “The capital of France is Paris”, accidentally unlearns “The capital of Spain is Madrid84% as strongly. (We unlearn only the tokens shown in purple.) It can even affect completely unrelated facts. Interestingly, incorrect facts are not disrupted. See Appendix A for more examples.
Similarly, unlearning biohazardous facts likely disrupts many benign biological concepts. This may explain why we can recover “unlearned” facts by retraining on unrelated biological text (deeb_unlearning_2024): retraining restores these disrupted benign concepts.
Figure 3: Disruption caused by unlearning a simple fact. We show how unlearning “The capital of France is Paris” disrupts the recall of other facts. We measure disruption using cosine similarity between the model’s update on the ”Paris” fact and the other evaluated fact. Activations column shows a slice of activations incoming into a middle layer MLP module at the token position right before the answer. Gradients column shows a slice of the gradients incoming into the same module during backpropagation when unlearning the answer (in purple).
In Figure 3, the activations, and to a lesser extent, the gradients, are very similar across different facts. This sheds light on why superficially similar facts are disrupted: most representations are not specific to the fact we are trying to unlearn, but more general. Since updates are computed from these “dirty” activations and gradients, other facts that share the same general representations are also affected. Therefore, preventing this requires a method to filter out those general representations.
@@ 3.4 Filtering out disruption is easier in activation and gradient space
A natural thing to try if we want to be selective is to limit which weights are updated. For example, sondej_robust_2025 showed unlearning improvements when allowing to modify only the weights where the signs of the unlearning and the retaining update are the same. Similarly, the A-GEM technique (chaudhry_efficient_2019) projects the weight updates to be orthogonal to the retaining updates to avoid performance disruption. Such projections have also been successfully used for unlearning (wu_unlearning_2025).
In Figure 4, the masked per weight row shows the effect of these filtering techniques. They significantly reduce the disruption (red), but some of it still escapes the filtering. That is because the control/retaining updates we use to decide which weights to filter out never match the actual disruption perfectly. (Compare the blue control pattern and the red disruption pattern.)
Figure 4: Comparison of two masking strategies. We show a slice of updates of a single weight matrix when unlearning “The capital of France is Paris”. Weights are colored green when an update successfully unlearns a paraphrased fact (”France’s capital is Paris”), red when it disrupts recall of a different fact (“The capital of Spain is Madrid”), and blue for a control fact disruption (“The capital of Italy is Rome”). Then we use the control fact disruption pattern to identify weights (or rows/columns) that are likely to be disruptive, and filter the unlearning update accordingly. Ideally we would want high unlearning transfer (green), with low disruption (red). Our approach of masking whole columns and rows removes disruption much more accurately.
Can we improve this filtering? Examining the update patterns in Figure 4 shows that both disruption and transfer appear as column- and row-wise stripes. Since weight updates are calculated as (activation × gradient) and thus are approximately low-rank, [^3] Strictly speaking their rank is equal to the number of tokens in the training batch, but most tokens have near-zero gradients, so the update could be approximated by a much lower-rank matrix. disruption is driven by certain rows and columns rather than isolated weights.
Since the disruption patterns shift within these columns and rows, it means that granular, per-weight filtering misses many harmful weights. Therefore, it is more effective to identify and remove whole faulty rows and columns (which is equivalent to ablating the corresponding dimensions in the activations and output gradients). Indeed, we see that doing so reduces the disruption-to-transfer ratio from 33% to 5%. Another advantage of intervening on whole columns and rows is reduced memory consumption: we operate on the activations and module output gradients (which are smaller) rather than the full weight updates.
@@ 4 Collapse of Irrelevant Representations
Following the findings from the previous section, we propose the activation and gradient-based method: Collapse of irrelavant representations (CIR).
@@
We have tried several ways to remove the representations which cause disruption. (By representations, we mean activations passed into model’s modules.) Simply ablating elements of the activations and gradients (as in Figure 4), while better than ablating individual weights, still struggles to cleanly remove the disruption. That is because representations exist in superposition (elhage2022superposition) – a single element can take part in encoding multiple representations, some relevant to the unlearning task, some not.
@@
We find that, instead of ablating, it is much more effective to project out the irrelevant representations. Defining them manually would be prohibitively tedious, so we approximate irrelevance by frequency: representations that are common across many training texts are likely irrelevant. Removing them leaves representations specific to the unlearned fact. Concretely, we locate the common subspace by centering the representation vectors (by subtracting the mean) and computing their principal components; we treat the mean as the “0th” PC, and when we say we collapse components we collapse the mean first. Equation 1 shows how to collapse activation PCs; we apply the same procedure to gradients.
@@
Based on grid searches shown in Figure 6(a) & 6(b) we chose to project 24 activation PCs and 36 gradient PCs. Performance plateaus for values between 12 and 48 (for both), making precise tuning unnecessary in this range. Removing the first few activation PCs is crucial.
@@
For each trained MLP module, we compute principal components (PCs) of its incoming activations and of the module-output gradients produced during backpropagation. Then, rather than using the usual update (activations × gradients), we first collapse the previously identified PCs, then compute the final weight update from the collapsed activations and gradients. PCs drift over time, so we recompute them periodically after every unlearning epoch. PCs may be estimated on any dataset, but we find best results when computing them on the unlearning corpus itself. This also makes the algorithm much more efficient, because we can reuse forward and backward passes for unlearning and for fetching activations and gradients.
We only intervene on MLPs, since this is where the model’s knowledge is stored (nanda2023factfinding). Also, collapsing representations on attention modules would be complex and specific to the model implementation. See Algorithm 1 for the pseudocode.
@@
CIR is compatible with any unlearning loss function and (optionally) any retain loss function. We first try loss functions which operate on the final logits, such as negative cross entropy, negative entropy (tamirisa_tamper-resistant_2024), or (proposed by us) directly minimizing the logit for the target token (but not below 0). We find the last one outperforms the prior loss functions, strongly preventing the model from recalling the harmful answer. However, it does not generalize to preventing recognizing the harmful answer in multiple-choice questions. While merely recognizing the answer is much less harmful, this result suggests that this approach may fail to generalize in other ways.
@@
In contrast, losses that target intermediate representations remove both recall and recognition of the harmful answers. The prior state-of-the-art representation breaking method is Circuit Breakers (zou_improving_2024), which minimizes (but only down to 0) the cosine similarity between current and initial activations of the residual stream.
We improve on this state-of-the-art in two ways. First, we note a problem with cosine similarity: it can be reduced not only by removing the original representation but also by adding a large, random direction. This can be disruptive, so we replace cosine similarity with the dot product. Indeed, in Figure 8 the dot-product loss disrupts the model much less for the same amount of unlearning, and we observe that cosine similarity-driven methods tend to grow activation norms.
Secondly, rather than breaking activations on the residual stream (which contains representations added by both MLPs and attention layers), we decided to work at the source and directly break the MLP outputs before they are added to the stream. Figure 6(c) shows that this improves unlearning-disruption tradeoff by 40%, and that targeting MLPs in layers 6–12 (for a 32-layer Llama 8B) is most effective. [^4] This also means we only need forward/backward passes on the first 12 layers, which is a major speedup. Therefore, our final unlearning loss is:
@@
We normalize by the average norm of the original MLP outputs so later layers (which have larger norms) do not dominate the loss. We also decide not to break representations at the <BOS> token position, since that would disrupt all texts, including benign ones. Finally, we train on a retain set with a representation-preserving loss that penalizes changes to the residual stream on retained data, i.e. $\text{resid\_stream}_{act}-\text{resid\_stream}_{orig\_act}||$, following the circuit breakers paper (zou_improving_2024).
Figure 5: WMDP-Cyber unlearning results. Circuit Breakers exhibit an abrupt unlearning reversal: the retain-loss term undoes earlier gains. A subsequent relearning run from the point of minimum accuracy proves even less robust. We also rerun CIR with a higher allowed disruption of 1% (baselines use 3%, but CIR’s high selectivity usually prevents reaching this threshold), but consistent with Section 3.1, unlearning gains are minimal. CIR with 0.1% allowed disruption already provides 30× higher unlearning robustness than the baselines.
5 Experimental setup